The toasters produced by a company have a normally distributed life span with a mean of 5.8 years and a standard deviation of 0.9 years, what warranty should be provided so that the company is replacing at most 4% of their toasters sold?
Answer by Boreal(15235)

In other words, we want the 4% ile of the distribution.
z(0.04)=-1.75
so -1.75=(x - man)/sd=(x-5.8)/0.9
-1.575=x-5.8
x=4.225 or 4.2 years